OpenStudy (anonymous):

3(x-2)^2+18=6 solve for x

OpenStudy (anonymous):

it has no real solutions

OpenStudy (anonymous):

why

OpenStudy (anonymous):

simplify, and you will find that the discriminant is less than zero

OpenStudy (anonymous):

oh, there's a simpler way, 3(x-2)^2=-12 (x-2)^2=-4 and a real number squared can't be negative

OpenStudy (anonymous):

it can if its imaginary

OpenStudy (anonymous):

yeah, and so it has no real solutions

OpenStudy (anonymous):

what would the negative number be

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

i dont know. lol